OpenStudy (jensine):

solve the equation 2x^2+28x+98=0

sam (.sam.):

2x^(2)+28x+98=0 Factor out the 2, 2(x^(2)+14x+49)=0 2(x+7)(x+7)=0 x+7=0 x=-7

OpenStudy (jensine):

thx u (: